A hypothetical population consists of eight individuals ages 13, 15, 17, 20, 22, 24, 28, and 30 years. (Enter your answers to three decimal places.)

(a) What is the probability that a person in this population is a teenager?


(b) What is the probability of selecting a participant older than 30?


(c) What is the probability of selecting a participant who is at least 20 years old?
Answer #1

a) Number of teenagers = 3

Hence,

P(Selecting teenager) = 3/8 = 0.375

b) P(Selecting participant older than 30) = 0

c) P(Selecting participant atleast 20 years old) = 5/8 = 0.625